# Rethinking Debiasing: Real-World Bias Analysis and Mitigation

### v3

@@ Abstract
Spurious correlations in training data significantly hinder the generalization capability of machine learning models when faced with distribution shifts in real-world scenarios. To tackle the problem, numerous debiasing approaches have been proposed and benchmarked on datasets intentionally designed with severe biases. However, it remains to be asked: 1. Do existing benchmarks really capture biases in the real world? 2. Can existing debiasing methods handle biases in the real world? To answer the questions, we revisit biased distributions in existing benchmarks and real-world datasets, and propose a fine-grained framework for analyzing dataset bias by disentangling it into the magnitude and prevalence of bias. We empirically and theoretically identify key characteristics of real-world biases poorly represented by existing benchmarks. We further introduce two novel biased distributions to bridge this gap, forming a systematic evaluation framework for real-world debiasing. With the evaluation framework, we focus on the practical setting of debiasing w/o bias supervision and find existing methods incapable of handling real-world biases. Through in-depth analysis, we propose a simple yet effective approach that can be easily applied to existing debiasing methods, named Debias in Destruction (DiD). Empirical results on real-world datasets in both image and language modalities demonstrate the superiority of DiD, improving the performance of existing methods on all types of biases.
@@ 1 Introduction
With the rapid development of machine learning, machine learning systems are increasingly deployed in high-stakes applications such as autonomous driving [1] and medical diagnosis [2], where incorrect decisions may cause severe consequences. As a result, the robustness to distribution shift is crucial in building trustworthy machine learning systems. One of the major reasons why machine learning models fail to generalize to shifted distributions in the real world [3, 4, 5] is because the existence of spurious correlation in training data [6]. Spurious correlation refers to the phenomenon that two distinct concepts are statistically correlated in the training distribution, yet uncorrelated in the test distribution for there is no causal relationship between them [7, 8]. For example, rock wall background may be correlated with the sport climbing in the training data, but they are not causally related and climbing can be indoors or on ice as well [9, 10, 11]. Furthermore, such spurious correlations within the data tend to be captured during training [12], resulting in a biased model that fails to generalize to shifted distributions. In this work, we refer to spurious correlation and bias in datasets interchangeably.
To tackle the problem, various debiasing methods [13, 12, 14, 9, 15, 16, 17, 18] have been proposed in recent years. And the effectiveness of the methods is benchmarked with synthetic [19, 12, 16] and semi-synthetic [9, 12, 17] (referred as "real-world dataset" in previous works) datasets designed to be severely biased. However, while these benchmarks are indeed biased, they are rough and lack thorough consideration of how data is truly biased in the real world. This raises two questions:
1. Do existing benchmarks really capture biases in the real world?
2. Can existing debiasing methods handle biases in the real world?
@@
To answer the first question, we revisit the biased distribution in existing benchmarks and real-world datasets, and propose a fine-grained framework for dataset bias analysis. Inspired by the framework proposed by Wiles et al. [6], which assumes the data is composed of some set of attributes, we further claim that analysis of dataset bias should be conducted on the more fine-grained feature (or value) level rather than attribute level, according to our observation on real-world biases. From the claim, we further propose our fine-grained framework that disentangles dataset bias into the magnitude of bias and the prevalence of bias, where the magnitude of bias generally measures how predictive (or biased) features are on the target task and the prevalence of bias generally measures how many samples in the data contain any biased feature. With our framework, we observe that the magnitude and prevalence of real-world biases are both low, in contrast with high magnitude and high prevalence biases in existing benchmarks. In section 3, we theoretically show that two strong assumptions are implicitly held by existing high bias prevalence benchmarks, which further validates our observation that real-world biases are low in bias prevalence.
As for the second question, due to the complexity of real-world biases, debiasing methods should be capable of handling various types of biased distributions rather than solely on the narrowed distribution in existing benchmarks. Thus, we introduce two new biased distributions inspired by real-world applications as a complement to the biased distribution in existing benchmarks, forming a systematic evaluation framework for real-world debiasing. We focus on debiasing methods w/o bias supervision[12, 9, 16, 15, 17, 18, 20], which is more practical as bias feature is expensive to annotate and sometimes even hard to notice [21]. Those methods generally involves a biased auxiliary model to capture the bias, along with techniques to learn a debiased model with the captured bias. (See the related work section in Appendix F for details) We condense them into a paradigm called debiasing with biased auxiliary model (DBAM). Our empirical results show that existing methods fail to handle real-world biases under various biased distributions.
We further conducted an in-depth analysis of existing methods and found that their effectiveness is reliant on the high bias prevalence of existing benchmarks and thus fails to handle real-world datasets with low bias prevalence. Finally, based on our analysis, we introduce a simple yet effective enhancement to existing methods. Experiments on real-world datasets in both image and language modalities show that our approach significantly boosts the capability of existing methods in handling real-world biases, improving their performances on both high and low bias prevalence datasets. To sum up, this work makes the following contributions:
@@
  • Identify key characteristics of real-world biases. We propose a fine-grained framework for analyzing bias in datasets. Based on the framework, we empirically and theoretically identified two key characteristics of real-world biases, previously overlooked. We then propose a systematic evaluation framework inspired by real-world applications for real-world debiasing.
  • Uncover critical failure mode of existing unsupervised debiasing methods. Evaluation of existing methods shows that they fail at handling real-world biases. We further prove that the failure is due to their reliance on unrealistic assumptions on the data distribution.
@@
  • Principled approach. A simple yet effective approach is proposed based on our analysis, which can be easily applied to existing unsupervised debiasing methods.
  • Empirical validation. Extensive empirical results on multiple real-world datasets, distributions, and modalities not only validate our analysis but also demonstrate the superiority of the proposed approach.
2 A Fine-grained Analysis on Bias in Datasets
@@
In this section, we first revisit the biases in existing debiasing benchmarks and biases in the real world. Then, we propose a new framework for assessing dataset bias. Based on the framework we show how existing benchmarks fail to represent bias in real-world conditions.
@@ 2.1 Revisiting spurious correlation in datasets
In the area of spurious correlation debiasing, multiple synthetic [19, 12, 16] and semi-synthetic datasets [9, 12, 17] have been adopted to benchmark the effectiveness of the debiasing methods. Generally, those synthetic datasets first select a target attribute as the learning objective [16], e.g. object, and another spurious attribute that could potentially cause the learned model to be biased, e.g. background. Then, certain sub-groups jointly defined by the target and spurious attributes, e.g. water birds with water background, are emphasized, i.e. synthesized or sampled from real-world datasets with much higher probability (usually above 95%) than the others in the biased dataset construction process, causing the corresponding spurious feature and target feature to be spuriously correlated, e.g. water background correlated with water bird [16]. Specifically, one such dominating subgroup is selected for every possible value of the spurious attribute, forming a "diagonal distribution", as shown in Figure 1(a) and 1(c). However, it is critical to examine whether this pattern truly aligns with the complexities of real-world biases.
@@ 2.2 Previous measures of spurious correlation
Background.
We assume a joint distribution of attributes $y^{1},y^{2},...,y^{K}$ with $y^{k}\in A^{k}$ where $A^{k}$ is a finite set. One of these $K$ attributes is the target of learning, denoted as $y^{t}$, and a spurious attribute $y^{s}$ with $t\neq s$. The definition of spurious correlation or the measure of bias magnitude is rather vague or flawed in previous works. We summarize the measures in previous works into three categories.
Target attribute conditioned probability.
Previous works [31, 19] measure spurious correlation according to the probability of a biased feature $a_{s}$ within the correlated class $a_{t}$: $Corr_{tcp}=P(y^{s}=a_{s}|y^{t}=a_{t})$. A higher value indicates a strong correlation.
Spurious attribute conditioned probability.
Some [23, 9, 32, 33] measure spurious correlation according to the probability of the correlated class $a_{t}$ within samples with biased feature $a_{s}$: $Corr_{scp}=P(y^{t}=a_{t}|y^{s}=a_{s})$. A higher value of the measure indicates a strong correlation.
Spurious attribute conditioned entropy.
Nam et al. [12] defined an entropy-based measure of bias. They use conditional entropy to measure how skewed the conditioned distribution is: $Corr_{sce}=H(y^{t}|y^{s})$, where $H$ is entropy. Values close to 0 indicate a strong correlation. This is an attribute-level measure, yet it is based on information theory.
@@
Spurious correlation should be measured at the feature level. As shown in Figure 1(a) and 1(c), the predictivity of every value in the spurious attribute is similar in existing benchmarks. However, this is not the case for real-world datasets, where it is clear that the predictivity of values in the spurious attribute varies greatly, as shown in Figure 1(b) and 1(d). Therefore, to deal with real-world biases, analysis of bias should be conducted on a more fine-grained value level, i.e. feature level, rather than attribute level in previous works [12]. Note that though $Corr_{tcp}$ and $Corr_{scp}$ are defined at the feature level, it is assumed by previous works [9, 19, 33] that it is consistent cross features in spurious attribute during benchmark construction, i.e. viewed as an attribute level measure.
The spurious attribute rather than the target attribute should be given as a condition. It is well recognized that the spurious attribute should be easier than the target attribute for the model to learn [12, 33]. Thus the spurious attribute should be more available to the model when learning its decision rules [33] and given as a condition when we define spurious correlation.
The marginal distribution of the target attribute should be accounted for. In $Corr_{tcp}$ and $Corr_{scp}$ measure of spurious correlation, the marginal distribution of the target attribute is not taken into account. This is inaccurate for even if the spurious and the target attribute are statistically independent, the value of $Corr_{tcp}$ and $Corr_{scp}$ could be high if the marginal distribution of spurious and target attribute is highly skewed, e.g. long-tail distributed [34, 35].
Diverge rather than predictivity should be used. While $Corr_{sce}$ satisfies the above requirements, it measures the entropy difference between the conditional and marginal distribution of the target attribute, i.e. the predictivity difference. This is still inaccurate for when the entropy of the distributions is the same, the conditional distribution could still be highly diverged from the marginal distribution, thus highly correlated with the spurious attribute. However, using divergence of the distributions accurately measures how the given condition affects the distribution shift of the target attribute.
@@ 2.5 Systematic evaluation framework for real-world debiasing
Based on our analysis, we further introduce two novel bias distributions inspired from real-world applications. Together with high magnitude high prevalence (HMHP) distribution in existing benchmarks [12, 16, 9, 17], we form a systematic evaluation framework for real-world debiasing.
@@
Note that datasets with low magnitude high prevalence (LMHP) bias do not exist due to the fact that high bias magnitude is the premise of high bias prevalence. Also, the bias distributions proposed are applicable to existing synthetic and semi-synthetic datasets mentioned in section 2. More details on how biased distributions are defined and how to synthesize datasets with given distributions can be found in Appendix B. In section 5, We further examine existing methods and our proposed method under real-world biases with the evaluation framework.
3 Theoretical Analysis
Data distribution.
Consider a classification task on binary target attribute $y^{t}\sim\{-1,+1\}$ and a binary spurious attribute $y^{s}\sim\{-1,+1\}$. Let the marginal distribution of the target and spurious attribute be $p^{t}_{+}=P(y^{t}=+1)$ and $p^{s}_{+}=P(y^{s}=+1)$. Then the joint distribution between $y^{t}$ and $y^{s}$ can be defined according to the conditional distribution of $y^{t}$ given $y^{s}=+1$, i.e. $\tau_{+}=P(y^{t}=+1|y^{s}=+1)$. We assume that feature $y^{s}=+1$ and $y^{s}=-1$ is correlated $y^{t}=+1$ and $y^{t}=-1$ respectively, i.e. $\tau_{+}>p^{t}_{+}$, in the following analysis.
$\rho_{+}=\tau_{+}-p^{t}_{+},\quad\rho_{-}=\tau_{-}-p^{t}_{-}$ (3)
Definition 2 (Biased Feature). We consider a feature $y^{s}=a$ biased if the ratio of its bias magnitude $\rho_{a}$ to its theoretical maximum $\rho^{max}_{a}=1-p_{a}^{t}$ is above certain threshold $0\leq\theta\leq 1$:
$\phi_{a}=\frac{\rho_{a}}{\rho^{max}_{a}}>\theta$
Definition 3 (High Bias Prevalence Distribution). We consider distribution as a high bias prevalence distribution only if both of the features in the spurious attribute are biased, i.e. $\phi_{+}>\theta,\phi_{-}>\theta$.
Proposition 1 (High bias prevalence distribution assumes matched marginal distributions). Assume feature $y^{s}=+1$ is biased. The high bias prevalence distribution, i.e. feature $y^{s}=-1$ is biased as well, implying that the marginal distribution of $y^{t}$ and $y^{s}$ is matched, i.e. $p^{t}_{+}=p^{s}_{+}$. Specifically, as $\theta$ approaches to $1$, the marginal distribution of $y^{s}$ approaches to that of $y^{t}$, i.e. $lim_{\theta\to 1}p^{s}_{+}=p^{t}_{+}$.
Proposition 2 (High bias prevalence distribution further assumes uniform marginal distributions even if they are matched). Given that the marginal distribution of $y^{s}$ and $y^{t}$ are matched and not uniform, i.e. $p=p^{s}_{+}=p^{t}_{+}<0.5$. The bias magnitude of sparse feature, i.e. $\rho^{*}_{+}$, is monotone decreasing at $p$, with $\lim_{p\to 0^{+}}\rho^{*}_{+}=-log(1-\phi_{+})$.The bias magnitude of the dense feature, i.e. $\rho^{*}_{-}$, is monotone increasing at $p$, with $\lim_{p\to 0^{+}}\rho^{*}_{-}=0$.
Remark 1. Proposition 2 reveals the fact that as the distribution of attributes becomes increasingly skewed, i.e. $p$ approaches 0, the magnitude of bias for features diverges, the magnitude of sparse increases while the magnitude of dense bias approaches 0. This results in biased sparse features and unbiased dense features, resulting in LP distributions.
4 Methodology
In this section, we dive into how real-world biases would raise a challenge to existing debiasing methods. Specifically, we first summarize and condense existing unsupervised debiasing methods into the debiasing with biased auxiliary model (DBAM) paradigm in section 4.1. Then we claim the insufficiency of existing DBAM paradigm when facing real-world biases in section 4.2. Finally, aiming at the insufficiency, we proposed our approach in section 4.3.
4.1 Debiasing with biased auxiliary model
In recent years, research in the field of debiasing has been more focused on the practical setting of debiasing w/o bias supervision. Though different in technical details, they generally adopt a biased auxiliary model to capture the bias, followed by techniques to learn a debiased model with the captured bias. The bias capture process is based on the assumption that the spurious attributes are easier and learned more preferentially than the target attribute, thus bias could be captured by an auxiliary model $M_{b}$ w/o bias supervision. To utilize $M_{b}$ for debiasing, sample re-weighing schemes have been widely incorporated into the training of the debiased model $M_{d}$, one common implementation of which is loss-based re-weighing $W(x)$ [12]:

### v4

# Towards Real-world Debiasing: Rethinking Evaluation, Challenge, and Solution

@@ Abstract
Spurious correlations in training data significantly hinder the generalization capability of machine learning models when faced with distribution shifts, leading to the proposition of numberous debiasing methods. However, it remains to be asked: Do existing benchmarks for debiasing really represent biases in the real world? Recent works attempt to address such concerns by sampling from real-world data (instead of synthesizing) according to some predefined biased distributions to ensure the realism of individual samples. However, the realism of the biased distribution is more critical yet challenging and underexplored due to the complexity of real-world bias distributions. To tackle the problem, we propose a fine-grained framework for analyzing biased distributions, based on which we empirically and theoretically identify key characteristics of biased distributions in the real world that are poorly represented by existing benchmarks. Towards applicable debiasing in the real world, we further introduce two novel real-world-inspired biases to bridge this gap and build a systematic evaluation framework for real-world debiasing, RDBench [^1] RDBench: Code to be released. Preliminary version in supplementary material for anonimized review.. Furthermore, focusing on the practical setting of debiasing w/o bias label, we find real-world biases pose a novel Sparse bias capturing challenge to the existing paradigm. We propose a simple yet effective approach named Debias in Destruction (DiD), to address the challenge, whose effectiveness is validated with extensive experiments on 8 datasets of various biased distributions.
@@ 1 Introduction
With the rapid development of machine learning, machine learning systems are increasingly deployed in high-stakes applications such as autonomous driving [1] and medical diagnosis [2], where incorrect decisions may cause severe consequences. As a result, the robustness to distribution shift is crucial in building trustworthy machine learning systems. One of the major reasons why machine learning models fail to generalize to shifted distributions in the real world [3, 4, 5] is because the existence of spurious correlation, i.e. biases, in training data [6]. Spurious correlation refers to the phenomenon that two distinct concepts are statistically correlated in the training distribution, yet uncorrelated in the test distribution for there is no causal relationship between them [7, 8]. For example, rock wall background may be correlated with the sport climbing in the training data, but they are not causally related and climbing can be indoors or on ice as well [9, 10, 11]. Furthermore, such spurious correlations within the data tend to be captured during training [12], resulting in a biased model that fails to generalize to shifted distributions. This lead to the proposition of various debiasing methods [13, 12, 14, 9, 15, 16, 17, 18, 19, 20] in recent years.
To benchmark the effectiveness of debiasing methods, both synthetic [21, 12, 16] and semi-synthetic [9, 12, 17] (referred to as "real-world dataset" in previous works) datasets with severe biases has been adopted as benchmarks. While the individual samples in semi-synthetic datasets are realistic as they are sampled from the real world rather than synthetic, both existing synthetic and semi-synthetic benchmarks follow some predefined biased distribution that lacks thorough consideration of how data is truly biased in the real world, as shown in Section 2.1. This raise the following question:
Does existing assumption on biased distributions align with the real world?
@@
This is a challenging question to answer given the complexity of biases in the real world and existing coarse-grained bias analysis measures (Section 2.2). Consequently, we first revisit the biased distribution in existing benchmarks and real-world datasets and propose a fine-grained framework for analyzing bias in datasets. Inspired by the framework proposed by Wiles et al. [6], which assumes the data is composed of some set of attributes, we further claim that analysis of dataset bias should be conducted on the more fine-grained feature (or value) level rather than attribute level, according to our observation on real-world biases. From the claim, we further propose our fine-grained framework that disentangles dataset bias into the magnitude of bias and the prevalence of bias, where the magnitude of bias generally measures how predictive (or biased) features are on the target task and the prevalence of bias generally measures how many samples in the data contain any biased feature. Empirical analysis on 8 real-world datasets across various modalities has shown that the magnitude and prevalence of real-world biases are both low, in contrast with high magnitude and high prevalence biases assumed by existing benchmarks. In section 3, we theoretically show that two strong assumptions are implicitly held by existing high bias prevalence benchmarks, which further validates our observation that real-world biases are low in bias prevalence.
Based on the empirical and theoretical insights on real-world biases, we introduce two novel type of biases inspired by real-world applications. Due to the complexity of real-world biases, debiasing methods should be capable of handling various types of biases and other real-world challenges, such as the multi-bias setting [22]. Thus, towards developing debiasing methods applicable in the real world, we propose a systematic evaluation framework for real-world debiasing that encompass various types of biases and settings to facilitate the debiasing field.
Furthermore, focusing on the setting of debiasing w/o bias label[12, 9, 16, 15, 17, 18, 23], which is more practical as bias feature is expensive to annotate and sometimes even hard to notice [24], we show that the proposed real-world biases pose a novel "Sparse bias capturing" challenge to the existing debiasing paradigm. Specifically, the sparse and scattered real-world biases make it difficult for existing methods to identify the unknown bias accurately, causing severe performance degradation. To tackle the challenge, we introduce a simple yet effective approach, named Debias in Destruction (DiD), that can be easily applied to existing methods. Extensive experiments on 8 datasets show that DiD significantly boosts the capability of existing methods in handling various types of biases. To sum up, this work makes the following contributions:
@@
  • Empirical and theoretical insights on biases in the real world. We propose a fine-grained framework for bias analysis. Based on the framework, we empirically and theoretically identified key characteristics of real-world biases, previously overlooked.
  • Systematic evaluation framework for real-world debiasing. Based on our insights, we further propose two novel real-world-inspired biases. Together with the multi-bias challenge and 8 existing benchmarks, we introduce a systematic evaluation framework for real-world debiasing to facilitate the development of real-world-applicable debiasing.
  • Uncover the "Sparse bias capturing" challenge in real-world debiasing. We show that the sparsity of real-world biases poses a unique challenge for debiasing w/o bias label.
@@
  • A simple-yet-effective approach to address the challenge. We propose an effective approach, which can be easily applied to existing debiasing methods. An extensive evaluation on 8 datasets of various distributions shows the effectiveness of the proposed approach.
2 A Fine-grained Empirical Analysis on Biased Distributions
@@
In this section, we first revisit the biased distributions in existing debiasing benchmarks and biases in the real world. Then, we propose a new framework for analyzing the biased distributions. Finally we propose our empirical findings on the consistent patterns in real-world biases.
@@ 2.1 Revisiting spurious correlation in datasets
In the area of spurious correlation debiasing, multiple synthetic [21, 12, 16] and semi-synthetic datasets [9, 12, 17] have been adopted to benchmark the effectiveness of the debiasing methods. Generally, those synthetic datasets first select a target attribute as the learning objective [16], e.g. object, and another spurious attribute that could potentially cause the learned model to be biased, e.g. background. Then, certain sub-groups jointly defined by the target and spurious attributes, e.g. water birds with water background, are emphasized, i.e. synthesized or sampled from real-world datasets with much higher probability (usually above 95%) than the others in the biased dataset construction process, causing the corresponding spurious feature and target feature to be spuriously correlated, e.g. water background correlated with water bird [16]. Specifically, one such dominating subgroup is selected for every possible value of the spurious attribute, forming a "diagonal distribution", as shown in Figure 1(a) and 1(c).
@@ 2.2 Previous measures of spurious correlation
Background. We assume a joint distribution of attributes $y^{1},y^{2},...,y^{K}$ with $y^{k}\in A^{k}$ where $A^{k}$ is a finite set. One of these $K$ attributes is the target of learning, denoted as $y^{t}$, and a spurious attribute $y^{s}$ with $t\neq s$. The definition of spurious correlation or the measure of bias magnitude is rather vague or flawed in previous works. We summarize the measures in previous works into three categories.
Target attribute conditioned probability. Previous works [34, 21] measure spurious correlation according to the probability of a biased feature $a_{s}$ within the correlated class $a_{t}$: $Corr_{tcp}=P(y^{s}=a_{s}|y^{t}=a_{t})$. A higher value indicates a strong correlation.
Spurious attribute conditioned probability. Some [26, 9, 35, 36] measure spurious correlation according to the probability of the correlated class $a_{t}$ within samples with biased feature $a_{s}$: $Corr_{scp}=P(y^{t}=a_{t}|y^{s}=a_{s})$. A higher value of the measure indicates a strong correlation.
Spurious attribute conditioned entropy. Nam et al. [12] defined an entropy-based measure of bias. They use conditional entropy to measure how skewed the conditioned distribution is: $Corr_{sce}=H(y^{t}|y^{s})$, where $H$ is entropy. Values close to 0 indicate a strong correlation. This is an attribute-level measure, yet it is based on information theory.
@@
Spurious correlation is better measured at the feature level. As shown in Figure 1(a) and 1(c), the predictivity of every value in the spurious attribute is similar in existing benchmarks. However, this is not the case for real-world datasets, where it is clear that the predictivity of values in the spurious attribute varies greatly, as shown in Figure 1(b) and 1(d). Therefore, to deal with real-world biases, analysis of bias should be conducted on a more fine-grained value level, i.e. feature level, rather than attribute level in previous works [12]. Note that though $Corr_{tcp}$ and $Corr_{scp}$ are defined at the feature level, it is assumed by previous works [9, 21, 36] that it is consistent cross features in spurious attribute during benchmark construction, i.e. viewed as an attribute level measure.
The spurious attribute rather than the target attribute is given as a condition. It is well recognized that the spurious attribute should be easier than the target attribute for the model to learn [12, 36]. Thus the spurious attribute should be more available to the model when learning its decision rules [36] and given as a condition when we define spurious correlation.
The marginal distribution of the target attribute should also be accounted for. In $Corr_{tcp}$ and $Corr_{scp}$ measure of spurious correlation, the marginal distribution of the target attribute is not taken into account. This is inaccurate for even if the spurious and the target attribute are statistically independent, the value of $Corr_{tcp}$ and $Corr_{scp}$ could be high if the marginal distribution of spurious and target attribute is highly skewed, e.g. long-tail distributed [37, 38].
It’s better to use diverge rather than predictivity. While $Corr_{sce}$ satisfies the above requirements, it measures the entropy difference between the conditional and marginal distribution of the target attribute, i.e. the predictivity difference. This is still inaccurate for when the entropy of the distributions is the same, the conditional distribution could still be highly diverged from the marginal distribution, thus highly correlated with the spurious attribute. However, using divergence of the distributions accurately measures how the given condition affects the distribution shift of the target attribute.
@@ 2.4 Observation on real-world biases
3 Theoretical Analysis on Biased Distributions
Data distribution.
Consider a classification task on the target attribute $y^{t}\sim\{a_{1}^{t},...a_{n}^{t}\}$ and a spurious attribute $y^{s}\sim\{a_{1}^{s},...a_{m}^{s}\}$. For any correlated target feature $a_{i}^{t}$ and spurious feature $a_{j}^{s}$, we have the marginal distribution of the target and spurious feature to be $p^{t}_{i}=P(y^{t}=a_{i}^{t})$ and $p^{s}_{j}=P(y^{s}=a_{j}^{s})$. Then the joint distribution between $y^{t}$ and $y^{s}$ can be defined according to the conditional distribution of $y^{t}$ given $y^{s}=a_{j}^{s}$, i.e. $\tau_{j}=P(y^{t}=a_{i}^{t}|y^{s}=a_{j}^{s})$.
$\rho_{j}=\tau_{j}-p^{t}_{i}$ (3)
Definition 2 (Biased Feature). We consider a feature $y^{s}=a_{j}^{s}$ biased if the ratio of its bias magnitude $\rho_{j}$ to its theoretical maximum $\rho^{max}_{j}=1-p_{i}^{t}$ is above certain threshold $0\leq\theta\leq 1$:
$\phi_{j}=\frac{\rho_{j}}{\rho^{max}_{j}}>\theta$ (4)
Definition 3 (High Bias Prevalence Distribution). We consider distribution as a high bias prevalence distribution only if both $y^{s}=a_{j}^{s}$ and $y^{s}\neq a_{j}^{s}$ are biased, i.e. $\phi_{j}>\theta,\phi_{\neq j}>\theta$.
Proposition 1 (High bias prevalence distribution assumes matched marginal distributions). Assume feature $y^{s}=a_{j}^{s}$ is biased. Then high bias prevalence distribution, i.e. feature $y^{s}\neq a_{j}^{s}$ is biased as well, implies that the marginal distribution of $a_{i}^{t}$ and $a_{j}^{s}$ is matched, i.e. $lim_{\theta\to 1}p^{s}_{j}=p^{t}_{i}$.
Proposition 2 (High bias prevalence distribution further assumes uniform marginal distributions even if they are matched). Given that the marginal distribution of $a_{j}^{s}$ and $a_{i}^{t}$ are matched and not uniform, i.e. $p=p^{s}_{i}=p^{t}_{j}<0.5$. The bias magnitude of sparse feature, i.e. $\rho^{*}_{j}$, is monotone decreasing at $p$, with $\lim_{p\to 0^{+}}\rho^{*}_{j}=-log(1-\phi_{j})$. The bias magnitude of the other features, i.e. $\rho^{*}_{\neq j}$, is monotone increasing at $p$, with $\lim_{p\to 0^{+}}\rho^{*}_{\neq j}=0$.
Remark 1. Proposition 2 reveals the fact that as the distribution of attributes becomes increasingly skewed, i.e. $p$ approaches 0, the magnitude of bias for features diverges, the magnitude of feature $a_{j}^{s}$ increases while the magnitude of other features $a_{\neq j}^{s}$ approaches 0. This results in extremely biased single feature and unbiased other features, resulting in LP distributions.
4 Methodology
In Section 4.1, we first propose a systematic evaluation framework for real-world debiasing based on our empirical (Section 2) and theoretical (Section 3) insights. Then, in Section 4.2 and 4.3, we dive into how real-world biases pose a challenge to existing methods. Finally, in Section 4.4, we propose a simple-yet-effective approach to adapt existing methods to the real-world scenarios.
@@ 4.1 Systematic evaluation framework for real-world debiasing
Based on our empirical and theoretical insights, we further introduce two novel types of bias inspired by real-world applications as follows. Together with high magnitude high prevalence (HMHP) distribution in existing benchmarks and other real-world challenges, we form a systematic evaluation framework for real-world debiasing. Please refer to Appendix B for full description of the framework.
@@
4.2 Existing paradigm for debiasing w/o bias label
In recent years, research in the field of debiasing has been more focused on the practical setting of debiasing w/o bias supervision. Though different in technical details, they generally adopt a biased auxiliary model to capture the bias, followed by techniques to learn a debiased model with the captured bias. The bias capture process is based on the assumption that the spurious attributes are easier and learned more preferentially than the target attribute, thus bias could be captured by an auxiliary model $M_{b}$ w/o bias labels. To utilize $M_{b}$ for debiasing, the generally shared heuristic is that BC samples should be relatively difficult for $M_{b}$ but not the debiased model $M_{d}$. One widely adopted implementation of the heuristic is the loss-based sample reweighing scheme $W(x)$ proposed by Nam et al. [12], which we use for our analysis: